# Highly Efficient Self-Adaptive Reward Shaping for Reinforcement Learning

### v3

@@ 4.2 Highly Efficient Beta Distribution Derivation
In this section, we present the methods to obtain the success and failure counts, $N_{S}(s_{i})$ and $N_{F}(s_{i})$, for deriving the Beta distributions. In tabular environments with low-dimensional, discrete, and finite state spaces, tracking these counts can be straightforwardly accomplished using two tables. However, most real-world environments involve high-dimensional, continuous, and infinite state spaces, making it impractical to maintain exact counts for every state.
To address this challenge, we use Kernel Density Estimation (KDE) to approximate the densities of successes and failures from accumulated experience. Specifically, we maintain two buffers, $\mathcal{D}_{S}$ and $\mathcal{D}_{F}$, to store the states in successful and failed trajectories, respectively. By treating these states as scattered data instances distributed across the state space, KDE estimates the density as:
@@
$\tilde{N}_{X}(\bm{s}_{i})=N\times\tilde{d}_{X}(\bm{s}_{i})=\frac{N}{|\mathcal{D}_{X}|}\sum_{j=1}^{|\mathcal{D}_{X}|}{\exp\left(-\frac{||\bm{s}_{i}-\bm{s}_{j}||^{2}}{2h^{2}}\right)},\quad X\in\{S,F\},$ (7)
where $h$ is a hyperparameter, the bandwidth of the Gaussian kernel.
We further integrate Random Fourier Features (RFF) (Rahimi and Recht,, 2007) to reduce computational complexity, as calculating the Gaussian kernel can be expensive, especially in scenarios involving high-dimensional state spaces and large buffers. RFF approximates the kernel function of the original $k$-dimensional states through an inner product of $M$-dimensional randomized features:
$K(\bm{s}_{i},\bm{s}_{j})\approx z(\bm{s}_{i})^{T}z(\bm{s}_{j}),\quad z(\bm{s})=\sqrt{\frac{2}{M}}\cos(\bm{W}^{T}\bm{s}+\bm{b}),$ (8)
where $z(\cdot)$ is the RFF mapping function with $\bm{W}=\big{[}\bm{w}^{(1)},\ldots,\bm{w}^{(M)}\big{]}\in\mathbb{R}^{k\times M}$ and $\bm{b}=\big{[}b^{(1)},\ldots,b^{(M)}\big{]}^{T}\in\mathbb{R}^{M}$ randomly sampled from the following spectral distributions:
@@
where $\bm{I}_{k}$ is the $k\times k$ identity matrix. Equation 9 is applied for the Gaussian kernel, while different kernels and the detailed derivations of the RFF method are provided in Appendix A.
@@ 4.2.1 Time and Space Complexity of SASR
Suppose the buffer size of $\mathcal{D}_{X}$ is $D$ and the batch size is $B$ per iteration, the computational complexity to compute the count $N_{X}$ is $O(MDB)$, indicating linear complexity (detailed derivation in Appendix B). RFF transforms nonlinear kernel computations into linear vector operations, significantly speeding up computation by leveraging the vectorization capabilities of GPUs (Dongarra et al.,, 2014). This demonstrates its higher efficiency compared to methods that rely on network updates and inferences involving extensive nonlinear computations. For an empirical comparison of inference time with and without RFF, as well as related discussions, please refer to Appendix C.
Given the retention rate $\phi$, the space complexity of maintaining two buffers is $O(\phi T|s|)$, where $T$ is the total iterations and $|s|$ is the size of a single state. However, storage space can be largely saved by leveraging the existing replay buffer in off-policy backbone RL algorithms, such as SAC (Haarnoja et al., 2018a, ) and TD3 (Fujimoto et al.,, 2018). By augmenting the replay buffer with a flag that marks each state as either success or failure, we can efficiently manage space requirements despite some additional indexing computations introduced during processing.
4.2.2 Implementation Details
Retention Rate. We introduce a hyperparameter, the retention rate $\phi\in(0,1]$, to regulate the volume and diversity of states stored in the buffers. Rather than storing all encountered states, we uniformly retain a specific portion of $\phi$. The motivation behind this are: (1) adjacent states in one trajectory tend to be highly similar, especially those near the initial state are repetitive and uninformative, retaining a portion of states can skip redundant states and increase sample diversity; (2) using a lower retention rate in the early stage keeps $N_{S}$ and $N_{F}$ lower, resulting in broader Beta distributions and preventing premature overconfidence.
@@ 4.3 The SASR Mechanism for RL agents
Building upon the SASR reward, we employ the soft actor-critic (SAC) algorithm by Haarnoja et al., 2018a as the foundation for our RL agent. Let $Q^{\mathcal{A}_{P}}_{\psi}$ be parameterized Q-network and $\pi^{\mathcal{A}_{P}}_{\theta}$ be parameterized policy network. The Q-network can be optimized by the following loss function:
where $Q_{{\psi}^{\prime}}$ is obtained from a secondary frozen target network to maintain a fixed objective (Mnih et al.,, 2015). It is worth noting that the environmental reward $r^{E}_{t}$ is retrieved from the replay buffer. Conversely, the shaped reward $R^{S}(s_{t})$ is computed in real-time using the most recently updated $N_{S}(s_{t})$ and $N_{F}(s_{t})$, thereby capturing the latest advancements in learning progress.
We optimize the policy network by maximizing the expected Q-value and the entropy of the policy $\mathcal{H}\big{(}\pi^{\mathcal{A}_{P}}(\cdot|s_{t})\big{)}$, following Haarnoja et al., 2018b :
@@
12:     $\tilde{N}_{S}=z(\bm{s}_{t})^{T}z(\bm{\mathcal{D}_{S}})$triangleright estimate the success counts
13:     $\tilde{N}_{F}=z(\bm{s}_{t})^{T}z(\bm{\mathcal{D}_{F}})$triangleright estimate the failure counts
@@
The flow of the SAC-embedded SASR algorithm is summarized in Algorithm 1. We maintain two buffers to store successful and failed states and estimate the success and failure counts from them. Subsequently, shaped rewards are computed by sampling from Beta distributions followed by a scale transformation. The shaped rewards, combined with environmental rewards, are used to update the agent’s Q-function and policy.
@@ 5 Experiments
We evaluate SASR in continuous control environments, following the guidelines established by Ma et al., 2024b , including seven MuJoCo tasks (Todorov et al.,, 2012), two robotic tasks (de Lazcano et al.,, 2023), and one physical simulation task (Towers et al.,, 2023), as shown in Figure 2. All tasks provide extremely sparse rewards, with a reward of $1$ granted only upon reaching the final objective within the maximum permitted steps. To ensure robust validation, we run $10$ instances for each setting using different random seeds and report the average results. We also maintain consistent hyperparameters and network architectures across all experiments, as detailed in Appendix E.
AntStand
AntSpeed
AntFar
AntVeryFar
WalkerKeep
HumanStand
HumanKeep
RobotReach
RobotPush
MountainCar
Figure 2: MuJoCo, robotic and physical simulation tasks with their objectives in our experiments. Detailed descriptions of each task are provided in Appendix G.
@@ 5.1 Comparison and Discussion
Baselines. We compare SASR with eight baselines to benchmark its performance: (a) RL with an Assistant Reward Agent (ReLara) (Ma et al., 2024b, ), (b) RL Optimizing Shaping Algorithm (ROSA) (Mguni et al.,, 2023), (c) Exploration-Guided RS (ExploRS) (Devidze et al.,, 2022), (d) Count-based static hashing exploration (#Explo) (Tang et al.,, 2017), (e) Random Network Distillation (RND) (Burda et al.,, 2018), (f) Soft Actor-Critic (SAC) (Haarnoja et al., 2018a, ), (g) Twin Delayed DDPG (TD3) (Fujimoto et al.,, 2018), and (h) Proximal Policy Optimization (PPO) (Schulman et al.,, 2017). Algorithms (a) to (e) are all reward shaping methods, involving either exploration bonuses or auxiliary agents to shape rewards, while algorithms (f) to (h) are advanced RL algorithms.
Figure 3 and Table 1 report the average episodic returns and standard errors throughout training. Our findings indicate that SASR surpasses the baselines in terms of sample efficiency, learning stability, and convergence speed. The primary challenge in these environments is the extremely sparse reward given after a long horizon, making exploration essential for timely obtaining successful trajectories. Although the exploration strategies of algorithms such as ExploRS, #Explo, and RND are designed to reward novel states, effectively expanding the early exploration space with the direct additional target, they continue focusing on discovering novel states, overlooking the implicit values of these states, which makes them fail to return to the final objectives.
@@
Table 1: The average episodic returns and standard errors of SASR and the baselines.
Algorithms AntStand AntSpeed AntFar AntVeryFar WalkerKeep HumanStand HumanKeep RobotReach RobotPush MountainCar
SASR 39.12$\pm$2.86 0.94$\pm$0.07 73.92$\pm$4.97 78.64$\pm$3.92 158.24$\pm$5.59 42.63$\pm$2.17 180.98$\pm$4.40 81.29$\pm$6.52 137.06$\pm$12.66 0.91$\pm$0.04
ReLara 28.66$\pm$1.82 0.33$\pm$0.02 67.77$\pm$4.30 64.07$\pm$4.17 77.14$\pm$8.77 29.72$\pm$1.85 160.31$\pm$7.30 103.56$\pm$7.18 58.71$\pm$6.98 0.89$\pm$0.01
ROSA 3.80$\pm$0.03 0.02$\pm$0.00 4.71$\pm$0.28 0.64$\pm$0.07 32.14$\pm$1.19 8.55$\pm$0.03 152.38$\pm$4.98 0.27$\pm$0.03 $0$0.90$\pm$0.02
ExploRS 4.52$\pm$0.04 0.01$\pm$0.00 5.42$\pm$0.22 1.67$\pm$0.10 2.47$\pm$0.13 8.63$\pm$0.03 158.09$\pm$4.42 0.79$\pm$0.04 0.20$\pm$0.08 -0.99$\pm$0.02
#Explo 6.79$\pm$0.50 0.00$\pm$0.00 12.40$\pm$1.66 0.55$\pm$0.09 131.56$\pm$5.40 28.73$\pm$1.79 160.60$\pm$7.04 4.19$\pm$0.42 6.31$\pm$0.85 0.79$\pm$0.02
RND 4.23$\pm$0.03 0.02$\pm$0.00 6.66$\pm$0.25 1.00$\pm$0.09 44.78$\pm$1.39 8.67$\pm$0.03 159.79$\pm$4.27 28.18$\pm$2.53 0.04$\pm$0.04 0.94$\pm$0.00
SAC 15.93$\pm$0.69 0.00$\pm$0.00 9.64$\pm$0.57 20.73$\pm$2.87 70.96$\pm$8.10 9.31$\pm$0.05 4.59$\pm$0.84 45.03$\pm$4.92 0.55$\pm$0.21 -0.05$\pm$0.02
TD3 0.00$\pm$0.00 0.00$\pm$0.00 0.81$\pm$0.02 0.81$\pm$0.02 18.62$\pm$0.75 5.72$\pm$0.04 0.55$\pm$0.03 0.00$\pm$0.00 0.00$\pm$0.00 0.00$\pm$0.00
PPO 4.54$\pm$0.04 0.00$\pm$0.00 6.33$\pm$0.24 1.80$\pm$0.11 33.77$\pm$1.11 8.36$\pm$0.03 138.13$\pm$12.64 79.52$\pm$10.80 0.00$\pm$0.00 0.93$\pm$0.00
@@
SASR outperforms the baselines primarily due to its self-adaptive reward evolution mechanism. In the early phases, SASR encourages exploration by injecting substantial random rewards to optimize the agent in multiple directions, thereby increasing the likelihood of collecting positive samples. Unlike novelty-driven exploration strategies that reward only novel states, SASR relaxes this assumption by rewarding random states, under the belief that all states are worth exploring, and this approach has proven empirically effective. Additionally, SASR leverages the success rate as auxiliary information to assess each state’s contribution to task completion, aligning with the agent’s primary objectives. As more data is collected, the success rate becomes more accurate, and the shaped reward provides more valuable guidance, significantly enhancing the exploitation and stabilizing the convergence toward optimal policies. Cooperatively, these strategies improve SASR’s sample efficiency and convergence stability in challenging tasks.
@@
While ReLara used a similar exploration mechanism to SASR by perturbing reward functions, it relies on an independent black-box agent, requiring more iterations to converge. In contrast, SASR’s success rate sampling is more direct, reducing the delay in converging to valuable information. ReLara’s advantage lies in incorporating the policy agent’s actions into reward construction, as seen in the RobotReach task, where the target point is randomly selected in each episode. In this case, ReLara outperforms SASR due to access to action information. This can also be achieved in SASR by treating actions as additional features in the state vector.

### v4

@@ 4.2 Highly Efficient Beta Distribution Derivation
In this section, we present how the success and failure counts, $N_{S}(s_{i})$ and $N_{F}(s_{i})$, are derived for the Beta distributions. To efficiently estimate these counts in high-dimensional, continuous, or infinite state spaces, we use Kernel Density Estimation (KDE) to approximate the densities of successes and failures from accumulated experience. Specifically, we maintain two buffers, $\mathcal{D}_{S}$ and $\mathcal{D}_{F}$, to store the states in successful and failed trajectories, respectively. By treating these states as scattered data instances distributed across the state space, KDE estimates the density as:
@@
$\tilde{N}_{X}(s_{i})=N\times\tilde{d}_{X}(s_{i})=\frac{N}{|\mathcal{D}_{X}|}\sum_{j=1}^{|\mathcal{D}_{X}|}{\exp\left(-\frac{||s_{i}-s_{j}||^{2}}{2h^{2}}\right)},\quad X\in\{S,F\},$ (7)
where hyperparameter $h$ is the bandwidth of the Gaussian kernel.
We further integrate Random Fourier Features (RFF) (Rahimi & Recht, 2007) to reduce computational complexity, as calculating the Gaussian kernel can be expensive, especially in scenarios involving high-dimensional state spaces and large buffers. RFF approximates the kernel function of the original $k$-dimensional states through an inner product of $M$-dimensional randomized features:
$K(s_{i},s_{j})\approx z(s_{i})^{T}z(s_{j}),\quad z(s)=\sqrt{\frac{2}{M}}\cos(\bm{W}^{T}s+\bm{b}),$ (8)
where $z(\cdot)$ is the RFF mapping function with $\bm{W}=\big{[}\bm{w}^{(1)},\ldots,\bm{w}^{(M)}\big{]}\in\mathbb{R}^{k\times M}$, and $\bm{b}=\big{[}b^{(1)},\ldots,b^{(M)}\big{]}^{T}\in\mathbb{R}^{M}$ is randomly sampled from the following distributions:
@@
where $\bm{I}_{k}$ is the $k\times k$ identity matrix. Equation 9 is applied for the Gaussian kernel, while different kernels and the detailed derivations of the RFF method are provided in Appendix A.1.
4.2.1 Implementation Details
Retention Rate. We introduce a hyperparameter, the retention rate $\phi\in(0,1]$, to regulate the volume and diversity of states stored in the buffers. Rather than storing all encountered states, we uniformly retain a specific portion of $\phi$. The motivations behind this are: (1) adjacent states in one trajectory tend to be highly similar, especially those near the initial state are repetitive and uninformative, retaining a portion of states can skip redundant states and increase sample diversity; (2) using a lower retention rate in the early stage keeps $N_{S}$ and $N_{F}$ lower, resulting in broader Beta distributions and preventing premature overconfidence.
@@ 4.2.2 Time and Space Complexity of SASR
Suppose the buffer size of $\mathcal{D}_{X}$ is $D$ and the batch size is $B$ per iteration, the computational complexity to compute the count $N_{X}$ is $O(MDB)$, indicating linear complexity (detailed in Appendix A.2). RFF converts nonlinear kernel computations into linear vector operations, significantly speeding up computation by leveraging the vectorization capabilities of GPUs (Dongarra et al., 2014). This highlights its superior efficiency compared to methods that rely on network updates and inferences involving extensive nonlinear computations.
Given the retention rate $\phi$, the space complexity of maintaining two buffers is $O(\phi T|s|)$, where $T$ is the total iterations and $|s|$ is the size of a single state. Moreover, storage space is significantly reduced by leveraging the existing replay buffer in off-policy RL algorithms, like SAC (Haarnoja et al., 2018a) and TD3 (Fujimoto et al., 2018). Specifically, we augment the replay buffer with a flag that marks each state as either a success ($\text{flag}=1$) or failure ($\text{flag}=0$). This allows for the efficient management of space requirements with minimal overhead from indexing.
For supporting experimental results in time and space complexity, please refer to Appendix A.3.
@@ 4.3 The SASR Mechanism for RL agents
12:     $\tilde{N}_{S}=z(s_{t})^{T}z(\mathcal{D}_{S})$triangleright estimate the success counts
13:     $\tilde{N}_{F}=z(s_{t})^{T}z(\mathcal{D}_{F})$triangleright estimate the failure counts
@@
Building upon the SASR reward, for demonstration, we employ the soft actor-critic (SAC) algorithm by Haarnoja et al. (2018a) as the foundational agent. Let $Q_{\psi}$ be the parameterized Q-network and $\pi_{\theta}$ be the parameterized policy network. The Q-network is optimized by the following loss function:
where $Q_{{\psi}^{\prime}}$ is obtained from a secondary frozen target network to maintain a fixed objective (Mnih et al., 2015). Notably, the environmental reward $r^{E}_{t}$ is retrieved from the replay buffer, conversely, the shaped reward $R^{S}(s_{t})$ is computed in real-time using the most recently updated $N_{S}(s_{t})$ and $N_{F}(s_{t})$, ensuring it reflects the latest learning progress.
We optimize the policy network by maximizing the expected Q-value and the policy entropy $\mathcal{H}\big{(}\pi_{\theta}(\cdot|s_{t})\big{)}$, following Haarnoja et al. (2018b):
The flow of the SAC-embedded SASR algorithm is summarized in Algorithm 1.
@@ 5 Experiments
We evaluate SASR in high-dimensional environments, including four MuJoCo tasks (Todorov et al., 2012), four robotic tasks (de Lazcano et al., 2023), five Atari games, including the well-known Montezuma’s Revenge Bellemare et al. (2013), and a physical simulation task (Towers et al., 2023), as shown in Figure 2. All tasks provide extremely sparse rewards, with a reward of $1$ granted only upon reaching the final objective within the maximum permitted steps. To ensure robust validation, we run $10$ instances per setting with different random seeds and report the average results. We also maintain consistent hyperparameters and network architectures across all tasks, detailed in Appendix A.6.
Figure 2: MuJoCo, robotic, Atari games and physical simulation tasks in our experiments. Detailed descriptions and the environmental reward models of each task are provided in Appendix A.7.
@@ 5.1 Comparison and Discussion
Baselines. We compare SASR with ten baselines to benchmark its performance: (a) the online Distributional Random Network Distillation (DRND) (Yang et al., 2024b), (b) RL with an Assistant Reward Agent (ReLara) (Ma et al., 2024a), (c) General Function Approximation Reward-Free Exploration (GFA-RFE) (Zhang et al., 2024), (d) RL Optimizing Shaping Algorithm (ROSA) (Mguni et al., 2023), (e) Exploration-Guided RS (ExploRS) (Devidze et al., 2022), (f) Count-based static hashing exploration (#Explo) (Tang et al., 2017), (g) Random Network Distillation (RND) (Burda et al., 2018), (h) Soft Actor-Critic (SAC) (Haarnoja et al., 2018a), (i) Twin Delayed DDPG (TD3) (Fujimoto et al., 2018), and (j) Proximal Policy Optimization (PPO) (Schulman et al., 2017). Algorithms (a) to (g) are all reward shaping methods, incorporating either exploration bonuses or auxiliary agents to shape rewards, while algorithms (h) to (j) are advanced RL algorithms.
Figure 3 shows the learning performance of SASR compared with the baselines, while Table 1 reports the average episodic returns with standard errors achieved by the final models over $100$ episodes. Our findings indicate that SASR surpasses the baselines in terms of sample efficiency, learning stability, and convergence speed. The primary challenge in these environments is the extremely sparse reward given after a long horizon, making exploration crucial for obtaining successful trajectories in a timely manner. Although exploration strategies of algorithms such as ExploRS, #Explo, and RND are designed to reward novel states, effectively expanding the early exploration space with direct additional targets, they continue prioritizing novelty, overlooking the implicit values of these states. As a result, they fail to return to the final objectives.
@@
Table 1: The average episodic returns and standard errors of all models tested over 100 episodes.
Tasks SASR DRND-online ReLara GFA-RFE ROSA ExploRS #Explo RND SAC TD3 PPO
AntStand 94.9$\pm$0.0 67.3$\pm$0.0 90.5$\pm$1.7 54.2$\pm$0.0 3.8$\pm$0.4 5.1$\pm$0.4 17.9$\pm$0.0 4.0$\pm$0.2 31.6$\pm$0.0 0.0$\pm$0.0 4.9$\pm$0.1
AntFar 139.8$\pm$0.0 93.2$\pm$0.0 115.7$\pm$0.0 86.4$\pm$0.0 1.0$\pm$0.0 12.0$\pm$4.2 75.1$\pm$0.0 4.6$\pm$1.6 25.3$\pm$0.0 1.0$\pm$0.0 7.8$\pm$0.0
HumanStand 79.8$\pm$2.0 50.6$\pm$0.0 76.2$\pm$0.7 58.2$\pm$0.0 8.8$\pm$0.0 9.3$\pm$0.0 72.7$\pm$0.0 9.3$\pm$0.1 9.9$\pm$0.0 5.5$\pm$0.0 9.0$\pm$0.1
HumanKeep 195.8$\pm$0.0 154.5$\pm$0.0 194.9$\pm$0.0 141.5$\pm$0.0 169.7$\pm$0.0 182.8$\pm$0.0 195.0$\pm$0.0 180.7$\pm$0.0 2.5$\pm$0.0 1.0$\pm$0.0 138.1$\pm$0.0
RobotReach 170.2$\pm$0.0 99.8$\pm$0.0 187.9$\pm$0.0 42.1$\pm$0.0 0.1$\pm$0.0 0.7$\pm$0.0 4.6$\pm$0.0 69.3$\pm$0.0 156.5$\pm$0.0 0.0$\pm$0.0 79.5$\pm$0.0
RobotSlide 132.3$\pm$1.3 127.2$\pm$0.0 111.6$\pm$2.0 115.8$\pm$2.0 11.2$\pm$0.9 4.3$\pm$0.1 3.5$\pm$0.0 4.8$\pm$0.2 0.7$\pm$0.2 0.5$\pm$0.4 0.2$\pm$0.2
RobotPush 167.1$\pm$0.0 122.2$\pm$0.0 166.9$\pm$0.0 49.1$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 3.7$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0
RobotPickPlace 1.0$\pm$0.0 1.0$\pm$0.0 1.0$\pm$0.0 0.5$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0
Pitfall 93.0$\pm$0.0 92.0$\pm$0.0 40.3$\pm$0.0 89.4$\pm$0.0 0.0$\pm$0.0 57.6$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 4.6$\pm$0.0 0.5$\pm$0.0 0.0$\pm$0.0
Frogger 14.2$\pm$0.0 11.7$\pm$0.0 11.6$\pm$0.0 7.9$\pm$0.0 9.8$\pm$0.0 8.3$\pm$0.0 11.9$\pm$0.0 10.5$\pm$0.0 0.8$\pm$0.0 0.7$\pm$0.0 0.0$\pm$0.0
Montezuma 6737.9$\pm$0.0 6828.5$\pm$0.0 2421.9$\pm$0.0 4755.3$\pm$0.0 4294.4$\pm$0.0 3971.5$\pm$0.0 1400.1$\pm$0.0 5494.3$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0
Solaris 42.1$\pm$0.0 21.3$\pm$0.7 20.3$\pm$0.0 26.3$\pm$0.0 0.1$\pm$0.0 17.0$\pm$0.0 1.2$\pm$0.8 9.8$\pm$0.0 6.0$\pm$0.0 0.4$\pm$0.0 1.5$\pm$0.0
Freeway 22.4$\pm$0.0 19.8$\pm$0.0 21.5$\pm$0.0 10.1$\pm$0.0 18.0$\pm$0.0 17.5$\pm$0.0 6.9$\pm$0.0 13.0$\pm$0.0 0.1$\pm$0.0 0.2$\pm$0.0 0.0$\pm$0.0
MountainCar 1.0$\pm$0.0 1.0$\pm$0.0 1.0$\pm$0.0 1.0$\pm$0.0 -0.9$\pm$0.0 -1.0$\pm$0.0 1.0$\pm$0.0 1.0$\pm$0.0 -0.1$\pm$0.0 0.0$\pm$0.0 0.9$\pm$0.0
@@
SASR outperforms the baselines primarily due to its self-adaptive reward evolution mechanism. In the early phases, SASR encourages exploration by injecting substantial random rewards, optimizing the agent in multiple directions, and increasing the likelihood of collecting positive samples. Moreover, since most states are initially classified as failures, their success rates decrease. As a result, unvisited states receive relatively higher rewards, encouraging further exploration. This mechanism resembles intrinsic motivation, assigning higher rewards to novel states, and effectively guiding the agent to expand the exploration space. As more data is collected, the success rate estimation becomes more accurate, the shaped reward provides more reliable guidance, enhancing exploitation and stabilizing convergence. Together, these strategies improve SASR’s sample efficiency and convergence stability in challenging tasks.
@@
While ReLara used a similar exploration mechanism by perturbing reward functions, it relies on an independent black-box agent, requiring more iterations to converge. In contrast, SASR’s success rate sampling is more direct, reducing delays in acquiring valuable information. ReLara’s advantage lies in incorporating the policy agent’s actions into reward construction, as seen in the RobotReach task, where the target point is randomly selected in each episode. In this case, ReLara outperforms SASR due to access to action information. However, SASR can achieve the same by incorporating actions as additional features in the state vector.